OpenStudy (anonymous):

Which value of a solves the equation? a^-2 = 1/64 A) -32 B) 4 C) -16 D) 8

OpenStudy (anonymous):

@alakara_kibo @kelliegirl33

OpenStudy (anonymous):

a^-2 = 1/64 1/a^2 = 1/8^2 a^2 = 8^2 now find a

OpenStudy (anonymous):

@rizwan_uet I don't know what a is.

OpenStudy (anonymous):

a is the parameter in your question

OpenStudy (anonymous):

@rizwan_uet Can't you just tell me?

OpenStudy (anonymous):

a = 8 is your answer

OpenStudy (anonymous):

well i was going to explain it but yes the answer is 8

OpenStudy (anonymous):

a^2 = 8^2 a^2 = 64 a = sq rt 64 a = 8